Manufacturer A has a profit margin of 2%, a total asset turnover of 1.8 and an equity multiplier of 5.1. Manufacturer B has a profit margin of 2.5%, a total asset turnover of 1.3 and an equity multiplier of 4.6. How much total asset turnover should Manufacturer B have to match Manufacturer As ROE 1.28 O3.19 O 2.20 O 1.6
Answer #1

Solution:

Manufacturer A's ROE = 2*1.8*5.1 = 18.36

Thus, required turnover = A's ROE / (B's Profit margin* Equity multiplier)

= 18.36 / (2.5*4.6) = 1.60

Thus, answer is option D) 1.60